Find two consecutive odd intergers such that the sum of the smaller and twice the larger is 79
Answer by checkley77(12569)

Let x & x+2 be the 2 integers.
x+2(x+2)=79
x+2x+4=79
3x=79-4
3x=75
x=75/3
x=25 answer for the smaller integer.
25+2=27 is the larger integer.
Proof:
25+2*27=79
25+54=79
79=79